OpenStudy (anonymous):

please help! due today, will fan if right! 5 questions. 1. What is the slope of the line? y=2x-7 slope=

OpenStudy (anonymous):

@iGreen ?

OpenStudy (anonymous):

It is already in Slope intercept form, so by looking at this, the slope should be 2.

OpenStudy (anonymous):

thanks! can you help me on some more questions?

OpenStudy (anonymous):

Absolutely:)

OpenStudy (anonymous):

what is the y-intercept of the line?? y= -2x-8

OpenStudy (anonymous):

In slope intercept form (y=mx+b) B= Y intercept. So, it will be -8.

OpenStudy (anonymous):

What is the slope, and what is the y-intercept of the line? y= 1/2x-5 slope= ?/? y-intercept=

OpenStudy (anonymous):

Slope is 1/2. Y-intercept is -5.

OpenStudy (anonymous):

What is the slope-intercept form of this equation? 6x + 4y = 8 A. y= 3/4x+2 B. y= -3/2x+4 C. y= -3/2x+2 D. y= 3/2x+4

OpenStudy (anonymous):

Should Be C.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

i got a 100% thank you!!!